OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10?

OpenStudy (anonymous):

a12 is 12th term a is 1st term 13=a+(12-1)*(-10) 13=a-110 a=123 S12= {1/2}*(a+a12) =0.5*(123+13) =0.5*136 =68

OpenStudy (anonymous):

1) 605 2) 660 3) 748 4) 816 these are the given answers

OpenStudy (anonymous):

ok so where he messed up was in the second part the equation is : S_n=n/29a_1+a_n So S_n=12/2(123+13) S_n=6(136) S_n=816

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

oh! that's what i put! great thank you!!

OpenStudy (anonymous):

no problem!